OpenStudy (anonymous):

A cone has a volume of 24 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of?

OpenStudy (campbell_st):

the volume of the cylinder is \[V_{cylinder} = \pi r^2h\] the cone with the same height and radius is \[V_{cone} = \frac{1}{3} \pi r^2 h\] using a little substitution you can find \[V_{cone} = \frac{1}{3} \times V_{cylinder}\] since you know the volume of the cone it should be easy to find the volume of the cylinder..